OpenStudy (anonymous):

Question 1: Find the volume of a cylinder that has a diameter of 8 cm and a height of 15 cm. Question 2: Find the surface area of a rectangular prism with length of 10 in, width of 6 in, and height of 5 in. Question 3: A rectangular prism has a length of 4", a width of 8", and a height of 12". If both the length and height double, how many times larger is the new volume than the original volume?

OpenStudy (anonymous):

What have you tried?

OpenStudy (anonymous):

1.) volume of a cylinder is -> pi r^2*h so pi 4^2*15=

OpenStudy (anonymous):

Soon as you ask a op if they've tried it, they leave lol.

OpenStudy (anonymous):

120

OpenStudy (anonymous):

@kique115

OpenStudy (anonymous):

that is the answer to 1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i am trying to find my old notes to help you with 3

OpenStudy (anonymous):

k thanks

OpenStudy (nikato):

for 1. recheck your calculation V=pi x r^2 x height =4^2 x 15xpi =?

OpenStudy (anonymous):

its 120

OpenStudy (nikato):

what's 16x15?

OpenStudy (anonymous):

240 @nikato

OpenStudy (nikato):

yes. 240 pi for 2. can you draw the net for the prism